# GenAINet: Enabling Wireless Collective Intelligence via Knowledge Transfer and Reasoning

### v2

@@ Abstract
GenAI and communication networks are expected to have groundbreaking synergies in 6G. Connecting generative artificial intelligence (GenAI) agents over a wireless network can potentially unleash the power of collective intelligence and pave the way for artificial general intelligence (AGI). However, current wireless networks are designed as a “data pipe” and are not suited to accommodate and leverage the power of GenAI. In this paper, we propose the GenAINet framework in which distributed GenAI agents communicate knowledge (high-level concepts or abstracts) to accomplish arbitrary tasks. We first provide a network architecture integrating GenAI capabilities to manage both network protocols and applications. Building on this, we investigate effective communication and reasoning problems by proposing a semantic-native GenAINet. Specifically, GenAI agents extract semantic concepts from multi-modal raw data, build a knowledgebase representing their semantic relations, which is retrieved by GenAI models for planning and reasoning. Under this paradigm, an agent can learn fast from other agents’ experience for making better decisions with efficient communications. Furthermore, we conduct two case studies where in wireless device query, we show that extracting and transferring knowledge can improve query accuracy with reduced communication; and in wireless power control, we show that distributed agents can improve decisions via collaborative reasoning. Finally, we address that developing a hierarchical semantic level Telecom world model is a key path towards a network of collective intelligence.
@@ I Introduction
To deploy LLMs on resource limited devices, various techniques have been proposed, such as model compression, inference acceleration, and parameter-efficient fine tuning. For instance, the T5-770M model has been shown to outperform the PaLM-540B in some tasks [3]. Moreover, paged attention proposed in vLLM can reduce the average inference time of LLMs down to 1/24 compared to vanilla methods [4]. Furthermore, QLoRA [5] is an efficient fine-tuning technique which can quantize an LLM to a 4-bit version with a small set of additional learnable low-rank adapters tuned using quantized weights. Although efficient in model size and inference time reductions, such techniques can degrade LLM’s performance, e.g., in terms of generalization capability, due to the non-modular nature of LLMs.
@@
To use LLMs in complex tasks solving, a multi-agent framework is required. An agent can observe the environment, plan a sequence of actions, create high-level reflections of experience in a memory stream, and formulate plans. For example, BabyAGI is a task-driven autonomous agent framework that can generate, execute and prioritize tasks in real-time. Auto-GPT can chain together LLM “thoughts” in an infinite loop of reasoning [6]. On the other hand, multi-agent systems are shown to largely enhance LLMs’ capability in task solving [7]. For example, CAMEL explores two role-playing communicative agents with autonomous cooperation [8]. In generative agents, believable human behaviors are simulated in a sandbox environment, where multiple agents interact in natural language to complete spontaneous tasks [9]. However, communication, computing, and storage efficiency are not considered in these frameworks, which are critical factors in wireless networks.
@@
Motivated by the reasons above, in this paper, we propose the GenAINet framework in which multiple GenAI agents communicate knowledge to perform effective reasoning and solve arbitrary tasks of particular applications, e.g., networks of vehicles, smart grid, and internet of things. In GenAINet, agents utilize LLMs to learn a high-level concept or abstract of data, plan reasoning path, and execute it to achieve a goal. The agents communicate knowledge to help each other’s planning and decision. In doing so, the performance can be improved with less communication and computing costs. The main contributions of this article are as follows:
  • A GenAINet architecture is discussed and possible approaches to integrate LLMs are proposed. We propose a unified agent architecture to manage both network protocols and applications;
  • A communication paradigm for GenAINet, built on semantic concepts and multi-path reasoning, is proposed. With semantic concepts, multi-modal raw data can be embedded on a common semantic space. With multi-step planning, we explain how an LLM can execute and optimize reasoning paths towards effective decisions. Semantic concepts learned from past reasoning and decision can be sent to or retrieved by remote agents, to improve independent or collaborative planning;
  • We investigate two case studies of wireless GenAINets. In the first case of question answering on a mobile device, we show transferring knowledge from a teacher to a student agent can improve query accuracy and communication efficiency; in the second case of wireless power control, we show that distributed GenAI agents can collaborate with reasoning to find a power allocation solution for a targeted data rate;
  • We discuss challenges and explore future research directions on GenAINet, including building a hierarchical world model with multi-level abstractions and planning to ground the model in real network scenario.
@@ II GenAI Network and Agent Architectures
Wireless GenAI agents can emulate human-like decision making process, providing a path towards sophisticated and adaptive GenAI networks. It can bring autonomy to network protocols and network applications. To achieve this, we propose prospect architectures for GenAI networks and GenAI agents, as illustrated in Fig. 1.
Figure 1: Proposed wireless GenAI network and agent architecture with protocol and application management.
II-A Wireless GenAI Network Architecture
@@
  • Distributed architecture: GenAI agents can be co-located with distributed NEs and manageme locally the data and control plans. The interaction between agents is self-organized, where agents exchange past decisions or knowledge to collaboratively accomplish the tasks.
@@
  • Independent architecture: Each application is controlled by individual GenAI agents. For example, in a vehicular network, the network agents control base stations, user equipments (UEs) and the traffic agents control cars, traffic lights.
@@
II-B Wireless GenAI Agent Architecture
To enable LLMs to interact with real-world scenarios, an agent has to be of a distributed structure, decision-wise and information-wise. An autonomous agent generally comprises four components: perception, action, planning, and memory. In what follows, we explain how these components can be designed to implement the proposed vision.
Perception is the part of an agent gathering useful and relevant information to an LLM. Possible information sources for perception include: 1) information sensed from the environment, e.g., channel state information (CSI), locations of user equipment (UE) and BS in a network, or on an autonomous car, the traffic density, speed and visual data of streets; 2) information received from other agents, such as messages reporting their state or responses to some requested information, or strategies learned from past decisions under different states. The information can be represented by using different modalities, from raw text, images, to abstracted graphs. The perceptor should fuse and encode raw data in a common embedding space, for the LLM to perform further decision-making.
Planning is the process of creating and optimizing actions over time to achieve a goal including: 1) sub-task decomposition, i.e., breaking down a high-level goal into sub-goals and actionable tasks; 2) synchronize and prioritize existing tasks between agents; 3) self-reflection, which evaluates and criticizes the past decisions to optimize the policy. Planning allows LLMs to solve complex tasks via multi-step reasoning, such as chain-of-thoughts (CoTs). Furthermore, such planning and reasoning can be conducted collaboratively among multiple agents by leveraging computing resources and tools on different devices efficiently.
Memory is the module for storing short-term experience and long-term knowledge for LLMs’ future planning and decisions. Short-term experience is the selected history of observations, actions, thoughts, and conclusions of agents. It is necessary for agents to efficiently reuse or adapt existing solutions or mechanisms instead of reasoning and extracting from scratch. Moreover, the long-term knowledge contains facts, methods, and contents which could be general or domain-specific depending on the role of an agent. Having such knowledge could effectively reduce the chance of an LLM to produce non-reliable or non-factual responses by using retrieval-augmented techniques. Possible knowledge representations and how LLMs leverage them to enhance the quality of inference will be covered in next section. Empowered by LLMs’ knowledge, GenAI agents can release the need for environment modeling or problem formulation, while finding the reasoning path towards task completion. This makes GenAI effective in handling complex, unseen problems.
@@ III Wireless GenAINet with Knowledge driven Communication and Reasoning
Existing LLMs and GenAI agents operate on raw data. For example, agents memorize and exchange information in natural language. Communication between AI models on the network edge is typically done by exchanging model weights, gradients, or hyper-parameters. However, such design schemes become inefficient when massive GenAI agents are connected. First, raw data contains a large amount of redundant information which is inefficient for communication; Second, optimizing LLMs over networks for different tasks is energy-consuming and may induce large latency. To address the aforementioned issues, we propose a semantic-native GenAINet. Empowered by LLM’s abstraction and planning capabilities, the agents communicate semantic concepts or abstracts learned from reasoning and decision making. In doing so, agents can learn to communicate or collaborate to achieve a goal, with reduced communication, computing and storage costs.
III-A Semantic knowledge representation
LLMs have been trained to understand semantics. Extracting knowledge from raw data to a concept space allows LLMs to effectively compose information for different tasks, which leads to better generalization. Knowledge includes information abstracted from a wide range of experiences. In GenAINet, this can be retrieved from common sources such as the web. Besides knowledge can be learned from an agent’s plan-and-execute experiences. For example, a car agent may learn to decelerate when observing other cars merging from the side lane. Another example, is when a solver agent may share with other solver agents an intermediate result about its calculation or computation process. Both retrieved and learned knowledge are thus vital for effective decision-making or generation.
Wireless GenAINet being constrained by communication, computing, and storage resources, knowledge transferred and stored in the network should satisfy minimality and sufficiency, that a minimal amount of information can solve a wide range of tasks or problems effectively. To realize this, we can decompose knowledge into two components: a) common knowledge in a domain which can be injected into LLMs via fine-tuning; b) evolving knowledge operating on top of common knowledge for specific tasks, which should be transferred in the network and used by LLMs. In GenAINet, managing common and evolving knowledge properly is the key to achieve high efficiency. In what follows, three possible representations for knowledge which are relevant for communication between LLMs are discussed.
@@ III-A Semantic knowledge representation
KG is a structured representation of relations between real objects or abstracted concepts. These relations can describe connections and causalities between entities in natural language, making it easily accessible for an LLM. Two vertices would be connected by an edge if they are correlated and the weight of the edge may indicate the degree of correlation. knowledge graph (KG) can assist an LLM to perform fast information retrieval, and efficient reasoning based on the structure of entities. Moreover, KG can also help LLMs to understand and generate more contextually relevant and coherent responses for complex tasks requiring factual responses.
@@
TE: Topological models, e.g., hypergraphs, simplicial complexes and cell complexes, can represent the intrinsic structure of data. They can model the low-order (vertices) and high-order (simplices, cells) relations between syntactic objects or abstracted concepts. Unlike conventional neural networks learning data structure on 1D sequence or 2D grid, topological embedding (TE) can exploit more flexible and dynamic structure. Moreover, the high-order topological structure can represent more complex and causal relations, such as different word combinations by semantic features. Therefore, TE can exploit more implicit latent structures than KG. The geometric features allows for flexible structure composition according to the changes of semantic features, which saves communication and computing resources.
III-B Multi-modal semantic reasoning
Semantic knowledge allows GenAI agents to compress, transfer, and retrieve new information. However, conventional structures built on the observation space lack generalization in new scenarios and domains. Therefore, building structures on a semantic concept space is essential. This requires aligning multi-modality raw data on a common embedding space and extracting semantic concepts. Techniques such as contrastive learning and multi-modal cross-attention (e.g., ImageBind [10]) have been used to align multi-modal data, extendable to RF signals as well. After training cross-modality encoders, semantic concepts can be extracted from raw data uniformly.
Semantic knoweldgebase can be built by learning a topological structure connecting semantic concepts. The first step is to exploit hidden semantic concepts from observations. Connections between semantic concepts can be learned from their semantic relations and causalities, which can be represented by different topological structures (e.g., edges and cliques). With logical combination of explicit concepts we can discover hidden concepts. This process can learn all possible semantic structures from raw data, then serving as a “world model” in the cloud. Subsequently, minimal and sufficient semantic concepts can be retrieved from semantic knoweldgebase to serve specific tasks, e.g., by clustering the concepts according to the semantic information required for a task.

### v3

@@ Abstract
genai and communication networks are expected to have groundbreaking synergies for 6G. Connecting Generative Artificial Intelligence (GenAI) agents via a wireless network can potentially unleash the power of Collective Intelligence (CI) and pave the way for Artificial General Intelligence (AGI). However, current wireless networks are designed as a “data pipe” and are not suited to accommodate and leverage the power of GenAI. In this paper, we propose the GenAINet framework in which distributed GenAI agents communicate knowledge (facts, experiences, and methods) to accomplish arbitrary tasks. We first propose an architecture for a single GenAI agent and then provide a network architecture integrating GenAI capabilities to manage both network protocols and applications. Building on this, we investigate effective communication and reasoning problems by proposing a semantic-native GenAINet. Specifically, GenAI agents extract semantics from heterogeneous raw data, build and maintain a knowledge model representing the semantic relationships among pieces of knowledge, which is retrieved by GenAI models for planning and reasoning. Under this paradigm, different levels of collaboration can be achieved flexibly depending on the complexity of targeted tasks. Furthermore, we conduct two case studies in which, through wireless device queries, we demonstrate that extracting, compressing and transferring common knowledge can improve query accuracy while reducing communication costs; and in the wireless power control problem, we show that distributed agents can complete general tasks independently through collaborative reasoning without predefined communication protocols. Finally, we discuss challenges and future research directions in applying Large Language Models in 6G networks.
Index Terms:
Collective intelligence, large language models, AI agents, semantic communications.
[…]
=-15pt
@@ I Introduction
[…]
Emerging Abilities refers to the capabilities of LLMs that appear suddenly and unpredictably as model size, computational power, and training data scale up. There are two typical emergent capabilities LLMs have been to able to perform: in-context learning [12] and instruct following [17, 18, 13]. In context learning is that LLMs are capable to perform unseen tasks provided with several demonstrative examples or the the description of the task in natural language. Such capability enable agents to accomplish different tasks in a varying environment without re-training or fine-tuning. Besides, LLMs can follow task instructions described in natural language without using explicit examples. These capabilities demonstrate the strong generalization ability of LLMs, making them suitable for diverse tasks in wireless networks.
Reasoning and Planning LLMs can perform logical or math reasoning to a certain level as showed in [19, 20]. For instance, LLMs can leverage Chain-of-Thoughts (CoTs) prompting, i.e., to obtain a final answer via multi-step intermediate reasoning state [14]. The capability of reasoning and planning are crucial to enable collective intelligence because: i) tasks in telecom context, e.g., Radio Resource Management (RRM) generally requires logical and math reasoning; ii) it is essential to decompose a high-level goal/intent to known executable tasks and plan them in a reasonable sequential order.
@@
Autonomous Cooperation LLMs can be further empowered into autonomous agents with external memory and tool utilization. An agent can observe the environment, plan a sequence of actions, create high-level reflections of experience in a memory stream, and formulate plans. For example, BabyAGI is a task-driven autonomous agent framework that can generate, execute and prioritize tasks in real-time [21]. Similarly, Auto-GPT can chain together LLM “thoughts” in an infinite loop of reasoning [22]. For solving even more complex tasks with LLMs, which are beyond the capability of a single agent, a multi-agent framework is required. On the one hand, multi-agent systems are shown to largely enhance LLMs’ capability in task solving [23]. For example, CAMEL explores two role-playing communicative agents with autonomous cooperation [15]. In generative agents, believable human behaviors are simulated in a sandbox environment, where multiple agents interact in natural language to complete spontaneous tasks [16]. On the other hand, adversarial interactions can also enhance overall performance through multi-agent debate as shown in [24, 25]. However, communication, computing, and storage efficiency are not considered in these frameworks, which are critical factors in wireless networks.
@@
Motivated by the advantages bringing by LLMs above and the drawbacks of the plug-in deployment of LLMs in a wireless network, in this paper, we propose the GenAINet  [^1] We adopt the terminology GenAI to refer to all generative foundation models including LLMs to maintain flexibility. framework. In our framework, agents leverage multi-modal LLMs to extract semantics in common embedding space from different data modalities. Then multiple GenAI agents exchange knowledge to perform effective reasoning and solve arbitrary tasks of particular applications, e.g., networks of vehicles, smart grid, and internet of things through effective semantic communication. The agents communicate knowledge to help each other’s planning and decision as well. In doing so, the performance can be improved with less communication and computing costs. The main contributions of this article are as follows:
  • We propose a novel GenAINet architecture that overcomes the drawback of existing frameworks by enabling dynamic task adaptation. Unlike current systems that are limited to fixed tasks, GenAINet integrates LLMs to handle evolving and arbitrary tasks arising from environmental changes or new user demands. This flexibility ensures the system remains adaptive and cost-efficient in dynamic scenarios.
  • We address the challenge of data heterogeneity by proposing a unified agent architecture capable of managing both network protocols and applications. This unified design allows for seamless integration of knowledge and decision-making mechanisms across diverse agents, eliminating the structural limitations of existing frameworks. Interaction between agents is no longer pre-defined; instead, reasoning and communication occur dynamically based on task requirements.
  • A communication paradigm for GenAINet is introduced, built on environmental semantics and multi-path reasoning. GenAINet mitigates the inefficiencies of raw data exchange by employing semantic-level communication, where semantics are extracted uniformly from multi-modal data using multi-modal LLMs. This reduces communication costs while retaining task-relevant information. Through multi-step planning, GenAINet leverages LLMs to optimize reasoning paths toward effective decisions, allowing experiences and mechanisms to be shared across agents to improve independent or collaborative planning.
  • Two case studies demonstrate GenAINet’s versatility in addressing limitations of traditional systems. In the first use case, semantic remote query on a mobile device, we show that transferring common knowledge from a teacher to a student agent improves query accuracy and reduces communication overhead through LLMs’ semantic compression. In the second use case, distributed wireless power control, we show that GenAI agents can solve power allocation problems independently, for arbitrary performance metrics, through collaborative reasoning, outperforming traditional fixed-policy approaches.
  • We explore challenges and future research directions, including deploying LLMs on edge devices, building hierarchical world models, and optimizing agent interactions to balance flexibility, scalability, and efficiency. These advancements position GenAINet as a transformative paradigm for addressing the limitations of existing wireless networks and AI approaches.
@@ II GenAI Network and Agent Architectures
Wireless GenAI agents can emulate human-like decision making process, providing a path towards sophisticated and adaptive GenAI networks. It can bring autonomy to network protocols and network applications. d in Fig. 1. To achieve this, we propose prospect architectures for GenAI agents, as shown in Fig. 1.
Figure 1: Proposed GenAINet architecture (left) and agent architecture (right) for protocol and application management. GenAINet leverages unified GenAI agents to manage communication protocols (e.g., PHY, MAC) and applications (e.g., robotics, vehicles). Each device is associated with an agent that communicates using semantic knowledge. The GenAI agent includes four components: perception (collects information from the environment and other agents), memory (stores long/short-term knowledge in formats like vector embeddings and knowledge graphs), action (available tools), and planning (reasoning and decision-making using LLMs).
II-A Wireless GenAI Agent Architecture
To enable LLMs to interact with real-world scenarios, an agent has to be of a distributed structure, decision-wise and information-wise. An autonomous agent generally comprises four components: perception, action, planning, and memory [26, 27]. In what follows, we explain how these components can be designed to implement the proposed vision and discuss their roles in wireless networks.
Perception is the part of an agent gathering useful and relevant information to an LLM. Possible information sources for perception include: 1) information sensed from the environment, e.g., channel state information (CSI), locations of User Equipment (UE) and Base Station (BS) in a network, or on an autonomous car, the traffic density, speed and visual data of streets; 2) information received from other agents, such as messages reporting their state or responses to some requested information, or strategies learned from past decisions under different states. The information can be represented by using different modalities, from raw text, images, to abstracted graphs. The perceptor should fuse and encode raw data in a common embedding space, for the LLM to perform further decision-making.
Planning is the process of creating and optimizing actions over time to achieve a goal including: 1) sub-task decomposition, i.e., breaking down a high-level goal into sub-goals and actionable tasks; 2) synchronize and prioritize existing tasks between agents; 3) self-reflection, which evaluates and criticizes the past decisions to optimize the policy. Planning allows LLMs to solve complex tasks via multi-step reasoning, such as Chain-of-Thoughts. Furthermore, such planning and reasoning can be conducted collaboratively among multiple agents by leveraging computing resources and tools on different devices efficiently.
Memory is the module for storing short-term experience (evolving knowledge) and long-term knowledge (common knowledge) for LLMs’ future planning and decisions. Short-term experience is the selected history of observations, actions, thoughts, and conclusions of agents. It is necessary for agents to efficiently reuse or adapt existing solutions or mechanisms instead of reasoning and extracting from scratch. Moreover, the long-term knowledge contains facts, methods, and contents which could be general or domain-specific depending on the role of an agent. Having such knowledge could effectively reduce the chance of an LLM to produce non-reliable or non-factual responses by using retrieval-augmented techniques. Possible knowledge representations and how LLMs leverage them to enhance the quality of inference will be covered in next section. Empowered by LLMs’ knowledge, GenAI agents can release the need for environment modeling or problem formulation, while finding the reasoning path towards task completion. This makes GenAI agents effective in handling complex, unseen problems.
II-B Wireless GenAI Network Architecture
@@
  • Distributed architecture: GenAI agents can be co-located with distributed NEs and manage locally the data and control planes. The interaction between agents is self-organized, where agents exchange past decisions or knowledge to collaboratively accomplish the tasks.
@@
  • Independent architecture: Each application is controlled by individual GenAI agents. For example, in a vehicular network, the network agents control BSs, User Equipments (UEs) and the traffic agents control cars, traffic lights.
@@
With GenAI agents introduced, the network management plane will be enhanced with automated policy generation and adaptation powered by LLM reasoning. The control plane will be enhanced with automated cooperation protocols between agents.
@@ III Wireless GenAINet with Knowledge driven Communication and Reasoning
Conventional communication systems operate on raw data. Moreover, agents memorize and exchange information from raw data regardless the relevance to their tasks. Besides, knowledge exchange between AI models on the network edge is typically done by exchanging model weights, gradients, or hyper-parameters, e.g., federated learning. However, such design schemes become inefficient when massive GenAI agents with gigantic model size are connected. First, raw data contains a large amount of redundant information which is inefficient for communication; Second, adapting LLMs over networks for different tasks is energy-consuming and may induce large latency. To address the aforementioned issues, we propose a semantic-native GenAINet. Pre-trained on massive amount of data, LLMs can understand the context of a problem. Therefore, it would be sufficient to merely exchange knowledge and information in brief/detailed textual description between GenAI agents to accomplish a goal collaboratively. Furthermore, wireless GenAINet being constrained by communication, computing, and storage resources, knowledge transferred and stored in the network should satisfy minimality and sufficiency, that a minimal amount of knowledge and information can solve a wide range of tasks or problems effectively. That said, how to efficiently exchange and store knowledge and information would be the key problem to enable wireless CI.
III-A Efficient Knowledge Representation
LLMs have been trained to understand semantics, the meaning of a message. This allows LLMs to effectively compose information for different tasks, leading to better generalization. Knowledge can be categorized into two main types: a) common knowledge in a domain including facts, known methods to solving problems, predefined standards and so on; b) evolving knowledge that builds on top of common knowledge for specific tasks, including useful experiences and effective mechanisms that LLMs have learned during operation. This highly abstracted knowledge should be efficiently shared among LLMs in a network. For example, common knowledge include predefined network parameter such as frequencies, bandwidth, and functions defined in Third Generation Partnership Project (3GPP) protocols and standards while evolving knowledge can be adaptive techniques for optimizing network performance in real-time scenarios, e.g., through Reinforcement Learning (RL). In what follows, three possible representations for knowledge which are relevant for communication between LLMs are discussed.
@@ III-A Efficient Knowledge Representation
kg is a structured representation of relations between real objects or abstracted concepts. These relations can describe connections and causalities between entities in natural language, making it easily accessible for an LLM. Two vertices would be connected by an edge if they are correlated and the weight of the edge may indicate the degree of correlation. Knowledge Graph (KG) can assist an LLM to perform fast information retrieval, and efficient reasoning based on the structure of entities. Moreover, KG can also help LLMs to understand and generate more contextually relevant and coherent responses for complex tasks requiring factual responses. To construct and maintain a KG tailored for 3GPP/IEEE standards and technical papers, key entities such as technical terms, specifications, and network components are extracted using NLP techniques like named entity recognition (NER). Relationships between these entities, such as dependencies or causal links, are identified through pattern matching and contextual analysis. The KG is then built by representing entities as vertices and relationships as weighted edges, with weights reflecting their relevance or significance. Continuous updates are essential as new standards and technical papers are published, ensuring the KG stays relevant and accurate. This dynamic KG enables GenAINet to perform efficient reasoning and information retrieval, supporting tasks like optimizing 5G network slicing or addressing emerging challenges in 6G networks.
@@
te: Topological models, e.g., hypergraphs, simplicial complexes and cell complexes, can represent the intrinsic structure of data. They can model the low-order (vertices) and high-order (simplices, cells) relations between syntactic objects or abstracted concepts. Unlike conventional neural networks learning data structure on 1D sequence or 2D grid, Topological Embedding (TE) can exploit more flexible and dynamic structure. Moreover, the high-order topological structure can represent more complex and causal relations, such as different word combinations by semantic features. Therefore, TE can exploit more implicit latent structures than KG. The geometric features allows for flexible structure composition according to the changes of semantic features, which saves communication and computing resources. Nevertheless, due to the complexity of topological models, the exploration of such structure is still on early stage. An example of applying TE in GenAINet can be seen in modeling dynamic network interactions in a 6G environment. For instance, a simplicial complex can represent communication relations among devices in a vehicular network, where vertices are individual vehicles, edges represent direct communication links, and higher-order simplices capture multi-device group interactions (e.g., cooperative sensing). By dynamically updating the simplicial complex based on semantic features like traffic patterns or interference levels, TE can efficiently model and adapt to changes in the network structure, enabling resource-efficient communication and computation while preserving critical causal relationships.
Vector embeddings, knowledge graphs (KG), and topological embeddings each offer distinct trade-offs in computational and communication complexity for knowledge representation. VE is computationally efficient, with low communication costs due to fixed-size embeddings, making it ideal for lightweight, real-time applications. However, it may lose granularity during compression. KG provides interpretable relationships and structured reasoning, with moderate computational complexity but higher communication costs for transmitting subgraphs. In contrast, TE captures rich, high-order interactions through complex topological structures, but this comes at the expense of very high computational and communication overhead, making it suitable only for specialized tasks. VE is most effective for efficiency-focused applications, KG excels in interpretability and reasoning, and TE is valuable for modeling intricate relational data in highly specialized scenarios.
To ensure effective integration, a robust mechanism for real-time updates and dynamic adaptation of the knowledge base (KB) is essential. Real-time updates can be achieved by enabling feedback loops between the querying device and the cloud-based KB. When new or ambiguous queries are encountered, the device can send feedback to the cloud, prompting the cloud LLM to refine or expand the KB with additional information. This dynamic process ensures that the KB evolves to accommodate emerging queries and remains aligned with the latest telecom standards and user needs. To balance data compression and query accuracy, the KB construction involves compressing semantic information into low-dimensional embeddings that capture essential knowledge while minimizing communication overhead. As we will show later in our first case study, the data compression rate of queries in different categories varies from one to others. Before compression, such knowledge learned during service could be provided to cloud LLM to adjust the compression rate carefully. This leads to a significant reduction of communication overhead especially for queries which are less sensitive to compression.
III-B Multi-modal Information Fusion and Reasoning
Suitable structure of knowledge representations allows GenAI agents to compress, transfer, and retrieve knowledge efficiently and timely. However, information of the environment, other agents’ states and actions obtained from observation space of different data modalities, e.g., images, audio, point could and Radio Frequency (RF) signals are generally correlated, thus introducing redundant information during transmission. Therefore, unifying the equivalent information from different data modalities or sources on a common semantic space is essential for efficient communication between GenAI agents. For the sake of simplicity, we select natural language as the semantic space for information in which LLMs effectively understand and manage. Meanwhile LLMs can serve as a bridge to connect the knowledge and semantics representing the perceived information. Techniques such as contrastive learning (e.g., CLIP [28]) and multi-modal cross-attention (e.g., ImageBind [29]) have been used to align multi-modal data, extendable to RF signals as well. After training cross-modality encoders, semantics can be extracted from raw data uniformly.